PLEASE SHOW THE COMPLET SOLUTION.
(Representation, Equation, Solution, Answer)
1. 8 years from now peter will be 26 years old. The current sum of the ages of peter and paul is 53. How old is paul right now?

Answer by tanjo3(60)

Let x be Peter's current age.

Then, x + 8 = 26. Solve for x.
x + 8 - 8 = 26 - 8
x = 18.
Peter is 18 years old right now.

Let y be Paul's current age.
Then y + 18 = 53. Solve for y.
y + 18 - 18 = 53 - 18
y = 35

Therefore, Paul is currently 35 years old.